OpenStudy (anonymous):

solve the following linear system x-2/3 + y+1/5 = 2 x+2/7 - y+5/7 = -2

OpenStudy (anonymous):

\[\frac{ x-2 }{ 3 } + \frac{ y+1 }{ 5 } = 2\] \[\frac{ x+2 }{ 7 } - \frac{ y+5 }{ 3 } = -2\]

OpenStudy (mathstudent55):

Multiply the first equation by 15 and the second eqaution by 21 to get rid of denominators.

OpenStudy (cwrw238):

first get rid of the fractions by multiplying each equation by the LCM of the 2 denominators which is 15 and 21 as mathstudent55 stated

OpenStudy (anonymous):

then what

OpenStudy (cwrw238):

then solve for x and y by the elimination or substitution method

OpenStudy (anonymous):

what is the equation then?

OpenStudy (anonymous):

show me the steps please

OpenStudy (cwrw238):

ok i'll convert the first one for you